What is limit of the relation between linear and angular accelaration..Precisely speaking where the relation can't be applied.

Grade:Upto college level

1 Answers

SAGAR SINGH - IIT DELHI
13 years ago

Dear student,

The angular acceleration can be defined as either:

{\alpha} = \frac{d{\omega}}{dt} = \frac{d^2{\theta}}{dt^2} , or
{\alpha} = \frac{a_T}{r}

Please feel free to ask your queries here. We are all IITians and here to help you in your IIT JEE preparation.

All the best.

Win exciting gifts by answering the questions on Discussion Forum. So help discuss any query on askiitians forum and become an Elite Expert League askiitian.

Now you score 5+15 POINTS by uploading your Pic and Downloading the Askiitians Toolbar  respectively : Click here to download the toolbar..

 

Askiitians Expert

Sagar Singh

B.Tech, IIT Delhi

sagarsingh24.iitd@gmail.com